### v1

@@
Reinforcement Learning with Verifiable Rewards (RLVR) has recently demonstrated notable success in enhancing the reasoning capabilities of large language models (LLMs), particularly in mathematics and programming tasks. It is widely believed that RLVR enables LLMs to continuously self-improve, thus acquiring novel reasoning abilities that exceed corresponding base models’ capacity. In this study, however, we critically re-examines this assumption by measuring the pass@k metric with large values of k to explore the reasoning capability boundary of the models across a wide range of model families, RL algorithms and math/coding benchmarks. Surprisingly, RLVR training does not, in fact, elicit fundamentally new reasoning patterns. We observed that while RL-trained models outperform their base models at smaller values of $k$ (e.g., $k$=1), base models can achieve a comparable or even higher pass@$k$ score compared to their RL counterparts at large $k$ values. Further analysis shows that the reasoning paths generated by RL-trained models are already included in the base models’ sampling distribution, suggesting that most reasoning abilities manifested in RL-trained models are already obtained by base models. RL training boosts the performance by biasing the model’s output distribution toward paths that are more likely to yield rewards, therefore sampling correct responses more efficiently. But this also limits their exploration capacity, resulting in a narrower reasoning capability boundary compared to base models. Similar results are observed in visual reasoning tasks trained with RLVR. Moreover, we find that, different from RLVR, distillation can genuinely introduce new knowledge into the model. These findings underscore a critical limitation of RLVR in advancing LLM reasoning abilities, which requires us to rethink the impact of RL training in reasoning LLMs and the need of a better training paradigm. Project Page: https://limit-of-RLVR.github.io
@@ 1 Introduction
The development of reasoning-centric large language models (LLMs), such as OpenAI-o1 Jaech et al. (2024), DeepSeek-R1 Guo et al. (2025), and Kimi-1.5 Team et al. (2025), has significantly advanced the frontier of LLM capabilities, particularly in solving complex logical tasks involving mathematics and programming. In contrast to traditional instruction-tuned approaches that rely on human-curated annotations Achiam et al. (2023); Grattafiori et al. (2024), the key driver behind this leap forward is large-scale Reinforcement Learning with Verifiable Rewards (RLVR) Lambert et al. (2024); Guo et al. (2025). RLVR starts with a pretrained base model or one fine-tuned on long chains of thought (CoT) data, optimizing it via reinforcement learning based on simple, automatically computable rewards. These rewards are determined by whether the model’s output matches a ground-truth solution in mathematics or passes unit tests in code, thus enabling scalability without human labeling. This framework has gained significant attention due to its simplicity and practical effectiveness. It is widely believed that RLVR enables models to autonomously incentivize advanced reasoning behaviors, such as enumeration, self-reflection, and iterative refinement, which are not present in base models Guo et al. (2025). Consequently, RLVR is viewed as a pathway to self-evolving LLMs with continuously expanding reasoning capabilities, potentially advancing us closer to stronger intelligence Guo et al. (2025).
Figure 1: (Left) The effect of RLVR on LLM’s reasoning ability. Search trees are generated by repeated sampling from the base and RLVR-trained models for a given problem. Grey indicates paths that are unlikely to be sampled by the model, while black indicates paths that are likely to be sampled. Green indicates correct paths, which has positive rewards. Our key finding is that all reasoning paths in the RLVR model are already present in the base model. For certain problems like Problem A, RLVR training biases the distribution toward rewarded paths, improving sampling efficiency. However, this comes at the cost of reduced scope of reasoning capacity: For other problems like Problem B, the base model contains the correct path, whereas that of the RLVR model does not. (Right) As RLVR training progresses, the average performance (i.e., pass@1) improves, but the coverage of solvable problems (i.e., pass@256) decreases, indicating a reduction in the model’s reasoning upper bound.
However, despite the empirical successes of RLVR, a critical question remains in the pursuit of continually self-evolving reasoning abilities:
Does RLVR really bring novel reasoning capabilities to LLMs? If so, what does the model learn from RLVR training?
@@
In order to rigorously answer this question, we first need to determine the boundary of reasoning capability of base and RL-trained models. Traditional metrics use single-pass success rates or average nucleus sampling Holtzman et al. (2019) to measure average-case performance. However, in our work, we highlight an important oversight behind such metrics: a model’s true reasoning potential may be underestimated if it fails on difficult problems after only a few attempts, even if it could succeed with more sampling. To address this, we employ a simple method, namely the pass@k metric Brown et al. (2024), where a problem is considered solved if any of the $k$ samples is correct. The idea is simple: What if we invest heavily in sampling for the base model with an enormous $k$? Can its performance match the RLVR-trained one? By giving the model a large number of trials to solve a problem, we are able to evaluate the boundary of reasoning capabilities of both base and RL-trained models. This provides a critical and rigorous test on whether the RLVR training yields fundamentally transcending capacity—enabling the model to solve problems that the base model cannot.
@@
We conduct extensive experiments across diverse benchmarks, including mathematics, code generation, and visual reasoning, spanning multiple LLM families, model sizes, and RL algorithms. Using the pass@k metric, we compare the performance of base models with their RL-trained counterparts. We find several perhaps surprising discoveries that may alter the conventional understanding of reasoning models and RLVR training, as follows:
@@
$\bullet$ RLVR-trained models perform worse than base models in pass@$k$ at large $k$ values. Although RL-trained models consistently outperform their base models in small $k$ as shown in previous work Guo et al. (2025), it is highly surprising that base models consistently surpass RL-trained models across all benchmarks and LLM families as $k$ increases. More notably, at sufficiently large $k$, base models achieve even higher pass@k scores than their RL-enhanced counterparts. This means that the base model without any RL training can already generate the correct answers to problems that were previously considered only solvable for RL-trained models by aggressively sampling different approaches. We manually check the CoTs of these correct answers and find most of problems has at least one CoT to be correct. This suggests that RL training does not improve, but even downgrade, LLMs’ potential scope of reasoning capability.
$\bullet$ RLVR boosts sampling efficiency but reduces the scope of reasoning capacity. To further investigate this phenomenon, we conduct perplexity analysis. Our findings show that reasoning paths generated by RLVR-trained models already exist within the base models’ output distribution with considerable probability density, indicating that these reasoning patterns and CoTs are not totally strange and unachievable for base models. RLVR training contributes to reasoning by biasing the model toward rewarded reasoning paths, thereby increasing the likelihood of sampling correct reasoning paths. However, this efficiency gain also comes at a cost: RL training reduces the model’s exploration capacity, resulting in smaller coverage of solvable problems at large $k$ (see  Figure 1 right), suggesting its reduced scope of reasoning capacity. This challenges the common belief that RLVR elicits transcending reasoning ability. Instead, the reasoning capacity boundary of RLVR-trained models may be bounded by the capabilities of base model. The effect of RLVR on LLM’s reasoning ability is illustrated in Figure 1 left.
@@
$\bullet$ RLVR algorithms perform similarly and remain far from optimal. Although different RL algorithms (e.g., PPO, GRPO, Reinforce++) show slight variations in performance, it does not make essential difference. As shown in Figure 7 top, we define the difference between the RL model’s pass@1 and the base model’s pass@$k$ (i.e., an potential performance upper-bound; we use $k=256$) as the sampling efficiency gap ($\Delta_{\text{SE}}$), which quantifies how effectively an RL algorithm approaches optimal sampling efficiency. We found that different RL algorithms produce only slight variations in $\Delta_{\text{SE}}$, yet the gap remains consistently large across all methods. This suggests that current RL approaches, which primarily work by improve sampling efficiency, are still far from optimal.
@@
$\bullet$ RLVR and distillation are fundamentally different. While RL improves sampling efficiency, distillation can genuinely introduce new knowledge into the model. As a result, distilled models often exhibit an expanded scope of reasoning capability beyond that of the base model by learning from distilled models, in contrast to RLVR-trained models whose capacity remains bounded by the base.
In conclusion, our results collectively indicate that RLVR, in its current form, is insufficient to incentivize transcending capabilities beyond the base model’s limits. Our work not only discovers findings that fundamentally challenge previous understanding on RLVR and its impact in reasoning models, but also implies that RLVR in itself might be inadequate to push forward the boundary of reasoning abilities.
Figure 2: Pass@$k$ curves of base models and their zero-RL-trained counterparts across multiple mathematical benchmarks. When $k$ is small, RL-trained models outperform their base versions. However, as $k$ increases to the tens or hundreds, base models consistently catch up with RL-trained models across all benchmarks and LLM families without exception. Eventually, base models surpass RL-trained models.
@@ 2 Preliminaries
In this section, we first introduce the fundamentals of RLVR. We then provide a detailed explanation of pass@$k$, the metric used to evaluate the reasoning boundary.
@@ 2.1 Reinforcement Learning with Verifiable Rewards
Verifiable Rewards. Let $\pi_{\theta}$ be an LLM with parameters $\theta$ that generates a token sequence $\mathbf{y}=(y_{1},\dots,y_{T})$ conditioned on a natural-language prompt $x$. A deterministic verifier $\mathcal{V}$ returns a binary reward: $r=\mathcal{V}(x,\mathbf{y})\in\{0,1\},$ where $r=1$ if and only if the model’s final answer is exactly correct (e.g., the numeric result matches the ground truth in math, or all unit tests pass in code). A format reward may also be added to encourage the model to explicitly separate the reasoning process from the final answer. The goal of RL is to learn a policy to maximize the expected reward: $J(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\cdot|x)}[\,r\,]\right]$ , where $\mathcal{D}$ is the distribution of prompts.
@@
where $r_{t}(\theta)=\pi_{\theta}(y_{t}|x,\mathbf{y}_{<t})/\pi_{\theta_{\text{old}}}(y_{t}|x,\mathbf{y}_{<t})$, and $A_{t}$ is the advantage estimated by a value network $V_{\phi}$. A KL divergence term is also applied, either as a reward penalty or directly in the loss function, to constrain the model from deviating too far from the original policy.
To reduce memory and computational overhead, several critic-free variants have been proposed. GRPO Shao et al. (2024) estimate the advantage with a normalized reward within a group of responses to the same question: $A_{i}=[r_{i}-\operatorname{mean}(\textbf{r})]/{\operatorname{std}(\textbf{r})}$, where $\textbf{r}=\{r_{1},\ldots,r_{G}\}$ denotes the set of rewards for a group of $G$ sampled responses. RLOO Ahmadian et al. (2024) instead adopts a leave-one-out baseline within each batch $\mathcal{B}$. Its adavantage is defined as $A_{i}=r_{i}-\tfrac{1}{|\mathcal{B}|-1}\sum_{j\neq i}r_{j}$.
Policy Gradient. PPO and its variants belong to the policy gradient class of reinforcement learning  Williams (1992); Sutton et al. (1998). These methods learn exclusively from on-policy samples, i.e., samples generated by the current LLM. In the context of verifiable binary rewards, the training objective is equivalent to maximizing the log-likelihood of samples with correct answers and minimizing the likelihood of those with incorrect answers.
Zero RL Training applies reinforcement learning directly to the base model without any supervised fine-tuning (SFT) on long CoT reasoning data Guo et al. (2025). To clearly study the effect of RLVR, we follow this zero-RL setting for all math tasks using open-source base models. However, for coding and visual reasoning tasks, no previous open-source work adopts a strict zero-RL setting, mainly due to training instability and limited effectiveness. Instead, these tasks typically use instruction-tuned models as starting points. We follow this convention, which is reasonable for studying the effect of RL on reasoning ability, as these models are primarily fine-tuned on general instruction-following data rather than long-form CoT reasoning data.
@@ 2.2 Metrics for LLM Reasoning Capacity Boundary
Accurately measuring the reasoning ability boundary of base and RL models is challenging, as methods like single-pass greedy decoding or the average of a few rounds of nucleus sampling Holtzman et al. (2019) reflect average-case performance rather than the model’s upper bound. For example, consider a traditional avg@8 approach, where 8 responses are sampled for an extremely difficult problem. If none of the responses are correct, the value is 0/8. However, after more trials such as 128, the model may solve the problem, indicating that the reasoning boundary can eventually overcome this challenge.

### v2

@@
Reinforcement Learning with Verifiable Rewards (RLVR) has recently demonstrated notable success in enhancing the reasoning performance of large language models (LLMs), particularly in mathematics and programming tasks. It is widely believed that, similar to how traditional RL helps agents to explore and learn new strategies, RLVR enables LLMs to continuously self-improve, thus acquiring novel reasoning abilities that exceed the capacity of the corresponding base models. In this study, we take a critical look at the current state of RLVR by systematically probing the reasoning capability boundaries of RLVR-trained LLMs across various model families, RL algorithms, and math/coding/visual reasoning benchmarks, using pass@k at large k values as the evaluation metric. While RLVR improves sampling efficiency towards correct paths, we surprisingly find that current training does not elicit fundamentally new reasoning patterns. We observe that while RLVR-trained models outperform their base models at smaller values of $k$ (e.g., $k$=1), base models achieve higher pass@$k$ score when $k$ is large. Moreover, we observe that the reasoning capability boundary of LLMs often narrows as RLVR training progresses. Further coverage and perplexity analysis shows that the reasoning paths generated by RLVR models are already included in the base models’ sampling distribution, suggesting that their reasoning abilities originate from and are bounded by the base model. From this perspective, treating the base model as an upper bound, our quantitative analysis shows that six popular RLVR algorithms perform similarly and remain far from optimal in fully leveraging the potential of the base model. In contrast, we find that distillation can introduce new reasoning patterns from the teacher and genuinely expand the model’s reasoning capabilities. Taken together, our findings suggest that current RLVR methods have not fully realized the potential of RL to elicit genuinely novel reasoning abilities in LLMs. This underscores the need for improved RL paradigms—such as continual scaling and multi-turn agent-environment interaction—to unlock this potential.    Project Page: https://limit-of-RLVR.github.io
Figure 1: (Left) The effect of current RLVR on LLM’s reasoning ability. Search trees are generated by repeated sampling from the base and RLVR-trained models for a given problem. Grey indicates paths that are unlikely to be sampled by the model, while black indicates paths that are likely to be sampled. Green indicates correct paths, which has positive rewards. Our key finding is that all reasoning paths in the RLVR model are already present in the base model. For certain problems like Problem A, RLVR training biases the distribution toward rewarded paths, improving sampling efficiency. However, this comes at the cost of reduced scope of reasoning capacity: For other problems like Problem B, the base model contains the correct path, whereas that of the RLVR model does not. (Right) As RLVR training progresses, the average performance (i.e., pass@1) improves, but the coverage of solvable problems (i.e., pass@256) decreases, indicating a reduction in LLM’s reasoning boundary.
@@ 1 Introduction
The development of reasoning-centric large language models (LLMs), such as OpenAI-o1 Jaech et al. (2024), DeepSeek-R1 Guo et al. (2025), and Kimi-1.5 Team et al. (2025), has significantly advanced the frontier of LLM capabilities, particularly in solving complex logical tasks involving mathematics and programming. In contrast to traditional instruction-tuned approaches that rely on human-curated annotations Achiam et al. (2023); Grattafiori et al. (2024), the key driver behind this leap forward is large-scale Reinforcement Learning with Verifiable Rewards (RLVR) Lambert et al. (2024); Guo et al. (2025). RLVR starts with a pretrained base model or one fine-tuned on long chains of thought (CoT) data, optimizing it via reinforcement learning based on simple, automatically computable rewards. These rewards are determined by whether the model’s output matches a ground-truth solution in mathematics or passes unit tests in code, thus enabling scalability without human labeling. This framework has gained significant attention due to its simplicity and practical effectiveness. In traditional RL settings such as game playing (e.g., Atari, Go), agents often autonomously discover new strategies and surpass even human-level performance through self-improvement Mnih et al. (2015); Silver et al. (2017). Inspired by this success, it is widely believed that RLVR similarly enables LLMs to autonomously develop novel reasoning patterns, including enumeration, self-reflection, and iterative refinement, surpassing the capabilities of their base models Guo et al. (2025). Consequently, RLVR has been considered a promising path toward continuously self-evolving LLMs, potentially bringing us closer to more powerful intelligence Guo et al. (2025).
However, despite its empirical success, the underlying effectiveness of current RLVR remains underexamined. This raises a fundamental question: Does current RLVR genuinely enable LLMs to acquire novel reasoning abilities—similar to how traditional RL discovers new strategies through exploration—or does it simply utilize reasoning patterns already in the base model?
@@
To rigorously answer this question, we must first assess the reasoning capability boundaries of both base and RLVR-trained models. Traditional evaluation metrics rely on average score from greedy decoding or nucleus sampling Holtzman et al. (2020), which reflects average-case behavior. However, these metrics risk underestimating the true potential of a model, especially if it fails on difficult problems after limited attempts, despite being capable of solving them with more sampling. To overcome this limitation, we adopt the pass@k metric Brown et al. (2024), where a problem is considered solved if any of the $k$ sampled outputs is correct. By allowing multiple attempts, pass@k reveals whether a model has the potential to solve a problem. The average pass@k across a dataset thus reflects the proportion of problems a model can potentially solve within $k$ trials, offering a more robust view of its reasoning boundary. This provides a rigorous test on whether the RLVR training yields fundamentally transcending capacity, enabling the model to solve problems that the base model cannot.
@@
Using the pass@k metric, we conduct extensive experiments across various benchmarks, covering multiple LLM families, model sizes, and RLVR algorithms to compare base models with their RLVR-trained counterparts. We uncover several surprising findings that offer a more comprehensive assessment of the effectiveness of current RLVR training and reveal the gap between existing RLVR methods and the ideal goals of RL-discovering genuinely new reasoning strategies:
@@
$\bullet$ Current RLVR models exhibit narrower reasoning coverage than their base models. In pass@$k$ curves, although RLVR models outperform their base models at small $k$, it is surprising that base models consistently surpass RLVR models across all benchmarks and LLM families as $k$ increases. This suggests that current RLVR training does not expand, and even reduce the scope of reasoning over solvable problems. Manual inspection of model responses shows that, for most problems, the base model can produce at least one correct CoT, implying that it can already generate correct reasoning paths for problems that were previously considered only solvable for RLVR models.
$\bullet$ Reasoning paths generated by current RLVR model already exist in its base model. To further investigate this phenomenon, we analyze the accuracy distribution. The results show that although RLVR improves average performance (i.e., pass@1) by sampling more efficiently on problems already solvable by the base model, it does not enable the model to solve new problems. Further perplexity analysis reveals that the reasoning paths produced by RLVR models already exist within the output distribution of the base model. These findings indicate that RLVR does not introduce fundamentally new reasoning capabilities and that the reasoning capacity of current RLVR models remains bounded by that of its base model. This effect of RLVR is illustrated in Figure 1 (left).
@@
$\bullet$ Current RLVR algorithms perform similarly and remain far from optimal. Treating the base model as an upper bound, we define the sampling efficiency gap ($\Delta_{\text{SE}}$), shown in Figure 8 (top), as the difference between an RL model’s pass@1 and the base model’s pass@$k$ (with $k=256$ as a proxy for upper-bound performance). This metric quantifies how closely an RL algorithm approaches the optimal bound. Across all algorithms (e.g., PPO, GRPO, Reinforce++), $\Delta_{\text{SE}}$ shows only minor variation yet remains consistently large, suggesting that current RLVR methods, while improving sampling efficiency, are still far from optimal.
@@
$\bullet$ RLVR and distillation are fundamentally different. While RLVR improves reasoning scores by more efficiently sampling high-reward outputs, it does not elicit new reasoning capabilities and remains constrained within the base model’s capacity. In contrast, distillation can transfer new reasoning patterns from a stronger teacher to the student. As a result, distilled models often demonstrate an expanded reasoning scope beyond that of the base model.
In conclusion, our findings show that current RLVR methods, while improving sampling efficiency, do not elicit novel reasoning beyond the base model’s capabilities. This highlights a gap between existing RLVR methods and the goals of reinforcement learning, underscoring the need for improved RL paradigms such as continual scaling, better exploration, and multi-turn agent interaction.
@@ 2 Preliminaries
In this section, we first outline the fundamentals of RLVR, then introduce the pass@$k$ metric to evaluate reasoning boundaries, and explain why it is preferred over alternatives like best-of-$N$.
Figure 2: Pass@$k$ curves of base models and their RLVR-trained counterparts across multiple mathematical benchmarks. When $k$ is small, RL-trained models outperform their base versions. However, as $k$ increases to the tens or hundreds, base models consistently catch up and surpass RL-trained models. More results on GSM8K and AMC23 can be found at Figure 9.
@@ 2.1 Reinforcement Learning with Verifiable Rewards
Verifiable Rewards. Let $\pi_{\theta}$ be an LLM with parameters $\theta$ that generates a token sequence $\mathbf{y}=(y_{1},\dots,y_{T})$ conditioned on a natural-language prompt $x$. A deterministic verifier $\mathcal{V}$ returns a binary reward: $r=\mathcal{V}(x,\mathbf{y})\in\{0,1\},$ where $r=1$ if and only if the model’s final answer is exactly correct. A format reward may also be added to encourage the model to explicitly separate the reasoning process from the final answer. The goal of RL is to learn a policy to maximize the expected reward: $J(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\cdot|x)}[\,r\,]\right]$ , where $\mathcal{D}$ is the distribution of prompts.
@@
where $r_{t}(\theta)=\pi_{\theta}(y_{t}|x,\mathbf{y}_{<t})/\pi_{\theta_{\text{old}}}(y_{t}|x,\mathbf{y}_{<t})$, and $A_{t}$ is the advantage estimated by a value network $V_{\phi}$. KL divergence term is optionally applied, to constrain the model from deviating too far from the original policy. More algorithms are introduced in Section D.4.
Policy Gradient. PPO and its variants belong to the policy gradient class of RL  Williams (1992); Sutton et al. (1998). These methods learn exclusively from on-policy samples, i.e., samples generated by the current LLM. In the context of verifiable rewards, the training objective generally maximizes the log-likelihood of samples with correct answers and minimizes the likelihood of those with incorrect answers.
Zero RL Training applies RL directly to the base model without any supervised fine-tuning (SFT) Guo et al. (2025). To clearly study the effect of RLVR, we follow this zero-RL setting for all math tasks using pretrained models as start model. However, for coding and visual reasoning tasks, open-source work typically uses instruction-tuned models as starting points, primarily due to the training instability and limited effectiveness of using a pure zero-RL setting. Following this convention, we compare the finetuned model with its RLVR-trained counterpart to focus solely on the effect of RLVR.